OpenStudy (anonymous):

Kelly tells you that when variables are in the denominator, the equation becomes unsolvable. "There is a value for x that makes the denominator zero, and you can't divide by zero," Kelly explains. Using complete sentences, demonstrate to Kelly how the equation is still solvable.

OpenStudy (zzr0ck3r):

let look at an example \(\frac{3}{x}=7\) for sure this does not even make sense if the denominator is \(0\), but that is really the only case we need to worry about. so we solve for x and get \(x=\frac{3}{7}\) and that is the solution and what do you know.... we dont need to even worry about the case where x is 0, because its not, its \(\frac{3}{7}\).

OpenStudy (zzr0ck3r):

now this question is worded badly, just because there is a variable in the denominator does not mean its solvable example \(\frac{x+1}{x}=1\) if we solve this we get \(x+1=x\) which implies \(1=0.....\color{red}{YIKES!!!!!!!!!!!!!!!}\)this shows that the equation is not solvable.